OpenStudy (anonymous):

WHAT IS THE DISTANCE BETWEEN POINTS M(6,-16) AND Z(-1,14), TO THE NEAREST TENTH

OpenStudy (saifoo.khan):

Use the formula.

OpenStudy (saifoo.khan):

here is the formula for that.

OpenStudy (anonymous):

CAN U JUST SHOW ME

OpenStudy (saifoo.khan):

SUre.

OpenStudy (saifoo.khan):

can u solve it

OpenStudy (anonymous):

ill tryyy it

OpenStudy (anonymous):

is it 289

OpenStudy (saifoo.khan):

Nope

OpenStudy (saifoo.khan):

\[\sqrt{(-1-6)^2+(14+16)^2}\]

OpenStudy (saifoo.khan):

\[\sqrt{(-7)^2+(30)^2}\]

OpenStudy (anonymous):

893

OpenStudy (saifoo.khan):

\[\sqrt{949}\]